OpenStudy (anonymous):

Which of the following is a solution of x2 + 5x = -2?

OpenStudy (mathstudent55):

If you have choices, plug in the values and see which ones work. If you don't have choices, use the quadratic formula or complete the square.

OpenStudy (anonymous):

5 plus or minus the square root of 33 divided by two. 5 plus or minus the square root of 17 divided by two negative 5 plus or minus the square root of 33 divided by two. negative 5 plus or minus the square root of 17 divided by two.

OpenStudy (anonymous):

\(x^2 + 5x = -2\\x^2 + 5x+2=0\\a=1, b=5, c=2\) solve using quadratic formula: \(\LARGE \begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & \\ \end{array}\)